# Position: Beyond Euclidean - Foundation Models Should Embrace Non-Euclidean Geometries

### v1

@@ 4 Towards Non-Euclidean Foundation Models
In the following roadmap on endowing foundation models with non-Euclidean geometry, we explore three progressive approaches: fine-tuning existing Euclidean models, building non-Euclidean models from scratch, and developing a hybrid framework combining both for optimal performance. We also discuss key implementation challenges of the roadmap, especially the computational efficiency of non-Euclidean operations.
@@ 4.1 Fine-tuning Existing Euclidean Foundation Models
The off-the-shelf Euclidean foundation model already encodes rich information from pretraining, offering a strong starting point for addressing the aforementioned challenges. A computationally efficient strategy is to adapt these models to non-Euclidean geometry, retaining their original capabilities while enabling generalization to data with non-Euclidean structures or relationships. We propose four strategies, shown in Figure 3(a): (1) Geometric prompt tuning; (2) Geometric low-rank adaptation; (3) Geometric knowledge distillation; and (4) Geometric transfer learning.
(1) Geometric Prompt Tuning. Prompt tuning offers a parameter-efficient alternative to full fine-tuning by introducing trainable, task-specific prompt tokens to the input. This approach helps mitigate catastrophic forgetting while requiring fewer trainable parameters. (Kirkpatrick et al., 2017; Han et al., 2024).
As there exist complex relations, such as hierarchical relationships between prefix prompts and downstream inputs, geometric prompts can be optimized through non-Euclidean spaces to align with the data geometry to better preserve structure and adapt to downstream tasks. For instance, trainable prompt and token embeddings could be introduced to better capture the topological relationships between prompts and text inputs.
(2) Geometric Low-Rank Adaptation. Low-Rank Adaptation (LoRA) offers an efficient way to adjust the model parameter space for downstream tasks given the observation of a low-rank structure (Hu et al., 2021). To equip the pre-trained model with non-Euclidean characteristics through geometric low-rank adaptation, we could operate the low-rank matrix multiplication directly on the manifold after projecting the input into non-Euclidean spaces, which avoids frequent switching and preserves the underlying geometric structure better (Yang et al., 2024a).
(3) Geometric Knowledge Distillation. Distilling knowledge into non-Euclidean spaces refers to transferring knowledge from a large, complex teacher model to a smaller, more efficient student model through geometric spaces. This approach utilizes the geometric properties of manifolds to teach the student to better inherit the manifold structure in the teacher model, such as by minimizing the gap between each layer’s output of both models, especially in high-dimensional spaces (Yang et al., 2022a; Hao et al., 2022) and resource-limited applications.
(4) Geometric Transfer Learning.
Geometric transfer learning aims to help foundation models learn across domains with aligned geometries, ensuring a much more effective and consistent knowledge transfer.
Geometry alignment objectives can be designed to supervise the transfer of geometric knowledge, such as hyperbolic contrastive learning for recommendation (Yang et al., 2024d; Ma et al., 2024), preserving the intrinsic structure of the target domain’s embedding space while retaining geometry-agnostic prior knowledge.
@@ 4.2 Pretraining from Scratch
Pretraining non-Euclidean foundation models requires addressing unique geometric challenges. We outline key components for adapting models to complex curvature-aware structures; see also Figure 3(b). A detailed mathematical formulation for each of the below sections
is presented in Table 4 in the Appendix.
Curvature Estimation. A data manifold’s curvature determines its intrinsic geometric properties, fundamentally affecting distance metrics, topological structure, and learning dynamics. Curvature estimation methods vary based on the data type. For graph data (e.g., networks, proteins), curvature is derived from topological properties, such as Ollivier-Ricci curvature or Gromov hyperbolicity (Ollivier, 2009; Jost & Liu, 2014; Yang et al., 2024a; Garcia Trillos & Weber, 2023). For non-graph data (e.g., text, images), curvature is typically estimated through learned embeddings (Khrulkov et al., 2020; Atigh et al., 2022) or techniques like Isomap (Tenenbaum et al., 2000) and UMAP (McInnes et al., 2018). We could also design learnable curvature within training pipelines using second-order statistics (Gao et al., 2022), reinforcement learning (Fu et al., 2021), and self-supervised learning (Sun et al., 2023, 2022).
Non-Euclidean Attention Mechanism. In non-Euclidean spaces, attention scores can be defined based on negative manifold distance $-d_{\mathcal{M}}(x,y)$ between queries and keys instead of dot products (Gulcehre et al., 2019; Shimizu et al., 2020; Chen et al., 2021b; Cho et al., 2023), with closer node pairs receiving higher attention weights. To aggregate attention, unified manifold centroids or tangent space operations can be used (Gu et al., 2019; Cho et al., 2023). Alternatively, linear attention mechanisms (Yang et al., 2024c) can be employed to improve computational efficiency by approximating traditional attention through unified tangent space operations.
Other Important Modules. Traditional Euclidean positional encoding, using sine and cosine functions (Vaswani et al., 2017), does not preserve the manifold structure in non-Euclidean spaces. Several approaches for non-Euclidean positional encoding include hyperbolic rotations (Yang et al., 2024c; Fein-Ashley et al., 2024), learned manifold encodings (Chen et al., 2021b), and direct embedding onto the manifold were proposed to represent token positions while maintaining geometric integrity. Residual connections should be formulated using isometric operations (He et al., 2025; van Spengler et al., 2023; Bdeir et al., 2024) to preserve geometric information across layers. Additionally, layer and batch normalization must be adapted to account for curvature (Yang et al., 2024c; Bdeir et al., 2024; van Spengler et al., 2023). One approach is to modify the softmax loss function and compute the probability distribution over tokens based on the manifold distance instead.
@@ 4.3 Hybrid Architectures
Hybrid architectures take a step further by merging both Euclidean and non-Euclidean foundation model architectures also provide a more universal inductive bias. We illustrate two promising strategies, also depicted in Figure 3(c).
Dynamic Geometry Adaptation Mechanisms. An intuitive way for hybrid modeling is to design a geometry-aware mechanism to shift dynamically between subspaces with different curvatures. The crux of this approach lies in how to perform efficient geometry adaption. A unified product manifold framework can serve as such a medium (Sun et al., 2024a), by utilizing a product layer to integrate diverse learnable curvature factors that adapt to different curvatures dynamically during the training process. Moreover, Mixture of Experts (MoE) (Zhou et al., 2022) provides a natural framework for such a hybrid paradigm using a geometry-aware sparse routing network aimed at minimizing average pairwise embedding distortion (Gu et al., 2019). This geometric routing selects the most appropriate Riemannian space considering data structure for the input (Guo et al., 2024), while avoiding interference from other spaces, thus naturally addressing issues of distortion and heterogeneity. Multi-Manifold Attention. To integrate more versatile underlying dependencies, we could transform self-attention into multi-manifold attention  (Konstantinidis et al., 2023; Kalitsios et al., 2024). Specifically, the input is embedded into a collection of manifolds (like Euclidean, hyperbolic, and spherical space) to represent differences in geometric structure across the data set. These geometric attention maps are then fused to produce a highly discriminative map for improved attention guidance.
4.4 Roadmap Implementation Challenges
Implementing and developing non-Euclidean foundation models come with unique challenges. One particular concern is that non-Euclidean operations are often complicated and require more computational resources, posing potential challenges when the scaling models to billions and trillions of parameters in practice. In particular, tangent-space-enabled methods and modules (Ganea et al., 2018; Chami et al., 2019; van Spengler et al., 2023), although intuitive by taking advantage of the Euclidean structure of the tangent spaces, incur significant computation overhead due to multiple mappings to and from the tangent bundle. In comparison, methods that operate directly on the manifold (Chen et al., 2021b; Yang et al., 2024c; He et al., 2025), while still computationally more expensive than Euclidean methods, typically have similar computational complexity as their Euclidean counterparts and offer significant speedup over tangent-space-based methods. This type of operation could be promising for managing the computational efficiency of non-Euclidean foundation models. Additionally, non-Euclidean models require fewer dimensions to embed complex structures when compared to Euclidean models, as seen in our discussion in Section 3.1. This enables the potential for non-Euclidean models to match the performance of Euclidean models with fewer parameters to build parameter-efficient foundation models. The reduction of parameter count could offset the computational overhead from non-Euclidean operations while offering additional benefits, such as the potential to continue the scaling law relationship between parameters and model performance.
Additional challenges include gaps in current research for more concrete analysis between embedding quality and model performance. Some previous works have explored how using manifolds that more accurately estimate the structure of the data can enhance performance in tasks related to GNNs and word embeddings (Gu et al., 2019). However, to the best of our knowledge, there currently lacks conclusive work directly connecting numerical distortion to downstream performance in foundation models, which is challenging as it could require prior knowledge of the ground-truth data geometry, compute resources to train multiple foundation models, and controlling and isolating the effects of distortion. Future works analyzing how distortion impacts the downstream tasks of foundation models would provide valuable insights to support better development of non-Euclidean methods for foundation models.
5 Alternative Views
While non-Euclidean geometries have clear theoretical benefits, there are practical challenges to consider. One of the main considerations is that operations such as parallel transport and exponential/logarithmic mappings add complexity, which might reduce some of the anticipated benefits in terms of efficiency. To tackle this, it is essential to develop specialized libraries that are optimized for these kinds of calculations, with efficient implementations of tensor operations that encode the underlying geometry.
@@ 5 Alternative Views
Another view is that as hardware capacities increase, simply scaling Euclidean models to higher dimensions might reduce distortion. However, as pointed out by Theorem 3.2 and Theorem 3.3, there is an upper bound in how much distortion reduction is possible. Previous works have also empirically shown that non-Euclidean models outperform Euclidean models even with scaled parameter counts, such as in the case of equivariant and non-equivariant models (Brehmer et al., 2024). Additionally, in many domains, such as molecular structures or rare languages, data scarcity results in brute-force scaling being ineffective. Non-Euclidean geometries, on the other hand, can capture important relationships even in lower-dimensional settings (Sarkar, 2011), making them efficient in both parameter usage and data requirements, offering better performance scalability with respect to model size and a more reliable option for domains with limited high-quality data.
@@ 6 Conclusion
Foundation models benefit from embracing non-Euclidean geometry to resolve the inherent mismatch between Euclidean geometry and the non-Euclidean nature of real-world data. Non-Euclidean geometries reduce distortion for complex structures and relationships while enabling efficient representations, which is critical for trillion-parameter scaling. Aligning architectures with data geometry can potentially mitigate hallucinations, boost efficiency, and unlock heterogeneous scaling. We encourage the community to prioritize three directions: unified curvature-adaptive foundation model frameworks, geometry-aware benchmarks, and studying manifold-emergent capability links. Embracing this paradigm will catalyze AI systems that better reflect the rich geometries of human knowledge and physical reality.
@@ A.1 Riemannian Geometry and Non-Euclidean Foundations
Hyperbolic Spaces. Hyperbolic spaces are Riemannian manifolds with constant negative curvature, i.e., with curvature $-K<0$. Common models for hyperbolic space include the Poincaré ball model $\mathbb{P}^{K,n}$ and the Lorentz hyperboloid $\mathbb{L}^{K,n}$, which have been extensively studied in the context of deep learning (Nickel & Kiela, 2017; Ganea et al., 2018). For points $\mathbf{x},\mathbf{y}\in\mathbb{L}^{K,n}$, their inner product $\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}$ is given by $\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=-x_{t}y_{t}+\mathbf{x}_{s}^{T}\mathbf{y}_{s}=\mathbf{x}^{T}\mathfrak{g}_{n}^{K}\mathbf{y}$ with $|\|\mathbf{x}\||_{\mathcal{L}}\coloneq\sqrt{|\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}|}$ being the Lorentzian norm. Formally, $\mathcal{L}^{n}$ is the point set $\mathcal{L}^{n}=\{\mathbf{x}\in\mathbb{R}^{n+1}:\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}=1/K,x_{t}>0\}.$ $\mathbb{P}^{n,K}$ is the $n$-dimensional sphere $S^{n}$ with radius $1/\sqrt{K}$, with the Riemannian metric $g_{x}^{\mathbb{P}}=\lambda_{x}^{2}g^{E}$, where $\lambda_{x}:=\frac{2}{1-c\|x\|^{2}}$ and $\mathfrak{g}^{E}$ is the Euclidean metric. Other models, such as the Klein model, also exist. These models are isometric, meaning that there is a smooth correspondence between points in different models that preserves distances, angles, and geodesics. This property allows for the selection of the most suitable model for a given application.

### v2

@@ 4 Towards Non-Euclidean Foundation Models
We propose a roadmap that explores three progressive approaches to incorporate non-Euclidean geometry in foundation models: fine-tuning existing Euclidean models, building non-Euclidean models from scratch, and developing a hybrid framework combining both for optimal performance. These methods could be evaluated with standard
@@ 4.1 Fine-tuning Existing Euclidean Foundation Models
Off-the-shelf pre-trained Euclidean foundation models are strong starting points as they already encode rich information. An efficient strategy is to adapt them to non-Euclidean spaces, thereby retaining their original capabilities and enabling generalization to data with non-Euclidean structures. We propose four strategies, shown in Figure 4(a): (1) Geometric prompt tuning; (2) Geometric low-rank adaptation; (3) Geometric knowledge distillation; and (4) Geometric transfer learning.
(1) Geometric Prompt Tuning. Prompt tuning offers a parameter-efficient alternative to full fine-tuning by introducing trainable, task-specific prompt tokens to the input, mitigating catastrophic forgetting while requiring fewer trainable parameters. [kirkpatrick2017overcoming, han2024parameter]. Geometric prompts can be optimized through non-Euclidean spaces to better align with the data geometry to and adapt to downstream tasks. For instance, trainable prompt and token embeddings could be introduced to better capture the topological relationships between prompts and text inputs.
(2) Geometric Low-Rank Adaptation. Low-Rank Adaptation (LoRA) offers an efficient way to adjust the model parameter space for downstream tasks [hu2021lora]. To equip the pre-trained model with non-Euclidean geometry through geometric low-rank adaptation, low-rank matrix multiplications could be performed directly on the manifold after projecting the input into non-Euclidean spaces, which better models the underlying geometric structure of the data [yang2024hyperbolic].
(3) Geometric Knowledge Distillation. Distilling knowledge into non-Euclidean spaces refers to transferring knowledge from a large, complex teacher model to a smaller, more efficient student model by utilizing manifold properties to teach the student to better inherit the teacher model’s geometric structure. An example is minimizing the gap between each layer’s output of both models, especially in high-dimensional spaces [yang2022geometric, hao2022learning] and resource-limited applications.
(4) Geometric Transfer Learning. Geometric transfer learning aims to help foundation models learn across domains with aligned geometries, ensuring a much more effective and consistent knowledge transfer. Geometry alignment objectives can be designed to supervise the transfer of geometric knowledge, such as hyperbolic contrastive learning for recommendation [10.1145/3627673.3679572, ma2024harec], preserving the intrinsic structure of the target domain while retaining geometry-agnostic prior knowledge.
@@ 4.2 Pretraining from Scratch
Pretraining non-Euclidean foundation models requires addressing unique challenges. We outline key components for adapting models to complex curvature-aware structures; see also Figure 4(b). A detailed mathematical formulation is presented in Table 4 in the Appendix.
Curvature Estimation. A manifold’s curvature determines its intrinsic geometric properties, such as distance metrics and learning dynamics. Curvature estimation methods vary based on data types. For graph data (e.g., networks, proteins), curvature can be derived from topological properties, such as Ollivier-Ricci curvature or Gromov hyperbolicity [ollivier2009ricci, jost2014ollivier, yang2024hyperbolic, trillos2023continuum]. For non-graph data (e.g., texts, images), curvature can be estimated from learned embeddings [khrulkov2020hyperbolic, atigh2022hyperbolic] or techniques like Isomap [tenenbaum2000isomap] and UMAP [mcinnes2018umap]. One could also design learnable curvature within training pipelines using second-order statistics [gao2022curvature], reinforcement learning [fu2021ace], and self-supervised learning [sun2023self, sun2022self].
Non-Euclidean Attention Mechanism. In non-Euclidean spaces, attention scores can be defined based on negative manifold distance $d_{\mathcal{M}}(x,y)$ between queries and keys instead of dot products [gulcehre2019hyperbolicAT, HNN++, chen2021fully, cho2023curve], with closer node pairs receiving higher attention weights. To aggregate attention, unified manifold centroids or tangent space operations can be used [gu2019mixedcurvature, cho2023curve]. Linear attention mechanisms [yang2024hypformer] can be employed to improve computational efficiency by approximating traditional attention through unified tangent space operations.
Other Important Modules. Traditional Euclidean positional encodings [vaswani2017attention, su2021roformer] do not preserve the manifold structure in non-Euclidean spaces. Several approaches for non-Euclidean positional encoding [chen2021fully, yang2024hypformer, fein-ashley2024hvt] were proposed to represent token positions while maintaining geometric integrity. Residual connections should be formulated using isometric operations [he2025lresnet, van2023poincar, Bdeir2024fully] to preserve geometric information across layers. Layer and batch normalization must also be adapted to account for curvature [yang2024hypformer, Bdeir2024fully, van2023poincar]. Loss function must also satisfy geometric constraints, such as computing the probability distribution over tokens based on the manifold distance instead.
@@ 4.3 Hybrid Architectures
Hybrid architectures take a step further by merging both Euclidean and non-Euclidean foundation model architectures to provide a more universal inductive bias. We illustrate two promising strategies, also depicted in Figure 4(c).
Dynamic Geometry Adaptation. An intuitive way for hybrid modeling is to design an efficient and geometry-aware mechanism that shifts dynamically between manifolds. Unified product manifold frameworks [10.1609/aaai.v38i8.28754] could enable layers to integrate diverse learnable curvature values that adapt to fine-grained geometric structures. Mixture of Experts (MoE) [zhou2022mixture] provides a natural framework for hybrid paradigms to use a geometry-aware sparse routing network by selecting the most appropriate geometry considering input structure [guo2024graphmoremitigatingtopologicalheterogeneity], addressing issues of distortion and heterogeneity.
Multi-Manifold Attention. Multi-manifold attention could lead to more versatile underlying dependencies  [konstantinidis2023multi, kalitsios2024dynamic], where the input is embedded into a collection of manifolds (including Euclidean) to represent differences in geometric structure across the dataset. These geometric attention maps are then fused to produce a highly discriminative map for improved attention guidance.
5 Alternative Views, Implementation Challenges, and Potential Solutions
Engineering Overhead. While non-Euclidean geometries have clear theoretical benefits, their operations can introduce significant engineering overhead that may offset efficiency gains. However, as mentioned earlier in the paper, non-Euclidean models require fewer dimensions to embed complex structures, as seen in our discussion in Section˜3.1 and 3.2. This enables the potential for non-Euclidean models to match the performance of Euclidean models with fewer parameters to offset the computational overhead while offering additional benefits, such as the potential to continue the scaling law relationship between parameters and model performance.
Nevertheless, designing efficient non-Euclidean operations is therefore a critical direction for future work. In particular, tangent-space-enabled methods [HNN, hgcn2019, van2023poincar] incur significant computation overhead due to multiple mappings to and from the tangent bundle. In comparison, methods that operate directly on the manifold [chen2021fully, yang2024hypformer, he2025lresnet], while still computationally more expensive than Euclidean methods, typically have similar computational complexity as their Euclidean counterparts. Thus, they could be promising for managing the computational efficiency of non-Euclidean foundation models. Additionally, it is essential to develop libraries, such as [he2025hypercore] that optimize these computations, with efficient implementations of tensor operations that encode the underlying geometry, geometric optimization techniques tailored to non-Euclidean spaces, and multi-node training architectures such as non-Euclidean flash attention to account for architectural differences as a result.
@@ 5 Alternative Views, Implementation Challenges, and Potential Solutions
Learning Inductive Bias with Euclidean Models. Another family of views is to let Euclidean foundation models learn geometric properties rather than hard-coding inductive biases. One argument holds that with increasing hardware capacity, scaling Euclidean models to higher dimensions could reduce distortion and capture geometry, as seen in CNNs and ViTs [dosovitskiy2020image]. A complementary argument is that improved data quality and quantity might allow Euclidean models to learn sufficient geometric structure. However, as shown in Theorem˜3.2 and 3.3, many desirable geometric biases cannot be faithfully learned by Euclidean foundation models. Previous works have also empirically shown that non-Euclidean models outperform Euclidean models even with scaled parameter counts, such as for equivariant and non-equivariant models [brehmer2024does]. Additionally, in many domains, such as molecular structures or rare languages, data scarcity results in brute-force scaling being ineffective. Non-Euclidean geometries, on the other hand, can capture important relationships even in lower-dimensional settings [sarkar2011low], making them efficient in data requirements, offering better performance scalability w.r.t. model size, and are more reliable for domains with limited high-quality data.
Distortion-Performance Connection. Additional views include gaps in analyzing the link between embedding quality and downstream foundation model performance. Prior work suggests that manifolds better capturing data structure can improve graph tasks and word embeddings [gu2019mixedcurvature]. Better semantic hierarchy modeling by incorporating non-Euclidean geometry could also lead to improved performance in LLMs and VLMs [he2025helm, desai2023hyperbolic]. However, to our knowledge, no conclusive studies connect distortion directly to downstream performance—a direction we advocate in this paper. This is challenging as it could require prior knowledge of the ground-truth data geometry, compute resources to train multiple foundation models, and isolating the effects of distortion. Future works in this aspect would provide valuable insights to support better development of non-Euclidean methods for foundation models. Additional analysis that could aid in this direction include analyzing embedding distribution in non-Euclidean models.
@@ 6 Conclusion
Foundation models benefit from embracing non-Euclidean geometry to resolve their inherent mismatch with the non-Euclidean nature of real-world data. Non-Euclidean geometries reduce distortion for embedding complex structures and relationships while enabling efficient representations, which is critical for trillion-parameter scaling. Aligning architectures with data geometry could mitigate hallucinations, boost efficiency, and unlock heterogeneous scaling. We encourage the community to consider three directions: unified curvature-adaptive foundation models, geometry-aware benchmarks, and studying manifold-emergent capability links. Embracing this paradigm will catalyze AI systems that better reflect the rich geometries of human knowledge and physical reality.
Acknowledgments
This work was supported in part by the National Science Foundation (NSF) IIS Div Of Information & Intelligent Systems 2403317. We also acknowledge support in part from the Silicon Valley Community Foundation, an Amazon research award, the Yale AI Engineering Research Grant from Yale Office of the Provost, and an LEAP-U Sponsored Research from Samsung Research America. Moreover, this research has greatly benefited from the discussions and research talks held at the IMS-NTU Joint Workshop on Applied Geometry for Data Sciences.
@@ A.1 Riemannian Geometry and Non-Euclidean Foundations
Hyperbolic Spaces. Hyperbolic spaces are Riemannian manifolds with constant negative curvature, i.e., with curvature K<0-K<0. Common models for hyperbolic space include the Poincaré ball model $\mathbb{P}^{K,n}$ and the Lorentz hyperboloid $\mathbb{L}^{K,n}$, which have been extensively studied in the context of deep learning [nickel2017poincare, HNN]. For points $\mathbf{x},\mathbf{y}\in\mathbb{L}^{K,n}$, their inner product $\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}$ is given by $\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=-x_{t}y_{t}+\mathbf{x}_{s}^{T}\mathbf{y}_{s}=\mathbf{x}^{T}\mathfrak{g}_{n}^{K}\mathbf{y}$ with $|\|\mathbf{x}\||_{\mathcal{L}}\coloneq\sqrt{|\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}|}$ being the Lorentzian norm. Formally, $\mathcal{L}^{n}$ is the point set $\mathcal{L}^{n}=\{\mathbf{x}\in\mathbb{R}^{n+1}:\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}=1/K,x_{t}>0\}.$ $\mathbb{P}^{n,K}$ is the $n$-dimensional sphere $S^{n}$ with radius $1/\sqrt{K}$, with the Riemannian metric $g_{x}^{\mathbb{P}}=\lambda_{x}^{2}g^{E}$, where $\lambda_{x}:=\frac{2}{1-c\|x\|^{2}}$ and $\mathfrak{g}^{E}$ is the Euclidean metric. Other models, such as the Klein model, also exist. These models are isometric, meaning that there is a smooth correspondence between points in different models that preserves distances, angles, and geodesics. This property allows for the selection of the most suitable model for a given application.
@@
Geometry of Embedding Spaces. Geometric assumptions are not only the result of neural network operations but are also inherently present when one chooses the latent space, whose underlying geometric properties determines not only how distances between representations are defined, but also what kind of operations are applicable in it. Consider a simple MLP example. In this model, each layer applies a linear transformation (e.g., $Ax+b$) followed by a non-linearity. Just by applying these transformations, one is implicitly assuming that the MLP’s representations lie in a geometrically Euclidean space that is flat and globally uniform, so that operations such as addition and scaling make sense everywhere in the same way. Simply put, one is implicitly assuming that the geometry of the space does not depend on location, and vectors can be freely added or transformed. In contrast, Riemannian manifolds are defined by a smooth manifold together with a Riemannian metric tensor. The smooth manifold defines the topological structure, which is not captured by the metric and is fundamentally not the same as just the set of $d$-dimensional real valued vectors (for example, for $d$-dimensional Lorentz hyperbolic space, $\mathcal{L}=\{x\in\mathbb{R}^{d+1}:\langle x,x\rangle_{\mathcal{L}}=1/K\}$), even though in reality one expresses the coordinates with real numbers. Thus, Riemannian manifolds have curvatures and do not, for example, support global linearity. This means that operations such as addition or scaling, which we take for granted in the standard MLP, cannot be directly applied on Riemannian manifolds. Instead, linear transformations must become location-dependent to account for curvature. Hence designing non-Euclidean neural networks typically requires redesigning the associated operations to conform to the underlying geometry.